# RiboFlow: Conditional De Novo RNA Sequence-Structure Co-Design via Synergistic Flow Matching

### v2

@@ 6 Conclusion
Impact Statement
Driven by the rapid advancement of artificial intelligence, the field of drug discovery is undergoing profound transformations. AI-based drug discovery offers the potential to significantly benefit humanity and enhance the expected quality of life. The remarkable success of mRNA vaccines during the COVID-19 pandemic, which saved countless lives, highlighted the importance of RNA engineering. Inspired by the broad application scenarios of RNA, we present RiboFlow, which, to the best of our knowledge, is the first deep generative model for the de novo design of RNA that targets specific molecules. We hope that RiboFlow can bring opportunities for the flourishing development of RNA therapeutic drugs and provide more research inspiration for pharmacologists and medical experts. Therefore, we hope that this important scientific contribution will receive the support of reviewers and the review committee.
Appendix A Dataset
A.1 Dataset Preparation
@@
(i) Data collection. We obtain raw structural data in CIF format from the RCSB PDB database using the advanced search query builder, which supports custom filtering based on specific structural attributes. The search is restricted to structures containing at least one RNA polymer chain and one ligand instance. These criteria are defined using the following query settings: Structure Attributes -> Number of Distinct RNA Entities >= 1 and Structure Attributes -> Total Number of Non-polymer Instances >= 1. This step serves as the initial data collection phase.
(ii) Ligand filtering. We adopt a portion of the ligand selection criteria from the HairBoss database (Panei et al., 2022), while relaxing certain rigid druggability restrictions, aiming to balance ligand diversity and effective identification. Specifically, the following rules are applied during our workflow: (1) the molecular weight of the ligand must fall within the range of 100-1000 Daltons, and (2) the ligand must not be a solvent molecule or a metal ion. Consequently, we expand the original ligand library beyond the 311 ligand classes used in HairBoss by incorporating an additional 237 ligand classes, resulting in a total of 548 unique ligand classes.
@@
(iv) Redundancy reduction. To reduce redundancy and improve dataset quality, we employ MMseqs2 (Steinegger & Söding, 2017) to cluster RNA sequences at a 90% sequence identity threshold. Notably, RNA sequences within the same cluster but bound to different ligands are still treated as distinct RNA-ligand pairs.
Figure S1: The collection pipeline of RiboBind dataset.
Figure S2: An example of RNA dynamic cropping. Based on a selected nucleotide, we randomly select a cropping length within the range of 30 to 200 to obtain the trimmed RNA-ligand pair for amplifying the existing data.
Through this comprehensive pipeline, we construct the RiboBind dataset, offering a substantially larger and more diverse collection of RNA-small molecule interactions compared to existing datasets. The final RiboBind dataset includes 1,591 RNA-ligand complexes and 3,012 RNA-ligand pairs. In comparison, as of December 2024, the publicly accessible HairBoss dataset comprises only 862 RNA-ligand complexes and 1,471 RNA-ligand pairs. This notable increase in both scale and diversity highlights the value of RiboBind as a robust resource for RNA-small molecule interaction studies.
A.2 Data Augmentation
@@
In order to ensure that the cropped RNA sequences retain interactions with small molecules, the cropping strategy is informed by interaction data from existing datasets. Specifically, we select RNA bases involved in interactions with the small molecules as candidate bases. By calculating the total interaction counts between each base and every atom in the small molecules, we can identify the top three bases with the highest interaction frequencies as cropping centers. The cropping length is randomly sampled within the range of 30–200 nucleotides. Using the selected base as the center, the RNA sequence is cropped symmetrically, with half the total length taken from both upstream and downstream of the base, forming the final RNA-ligand pair. This approach ensures that the cropped RNA sequences retain critical interaction information for downstream tasks.
@@
Appendix B More Results and Analysis
B.1 The Impact of Pre-training Steps on RNA Generation
To fully demonstrate the impact of different pre-training steps on RNA generation, we comprehensively show the changes in model performance under various pre-training steps. We additionally introduced the scRMSD metric to evaluate the performance changes. During the pre-training stage, a batch size of 32 is selected, utilizing 4 A100 80GB accelerator cards, completing 200K pre-training steps in nearly 20 hours. The experimental results are shown in Table S1.
Table S1: Detailed results of unconditional generation under different pre-training steps.
@@ B.1 The Impact of Pre-training Steps on RNA Generation
Figure S3: Comparison of probability distribution histograms of nucleotide bond distances, bond angles, and torsion angles between the generated RNA and the real RNA structure.
B.2 The Local Structure of RNA Analysis
@@
To demonstrate the validity of the generated RNA structures, we conduct a detailed comparison between the RNA structures generated by RiboFlow-Codesign within the [50,150] range and the actual RNA structure distributions. We primarily illustrate histograms of the probability distributions for the bond distances between nucleotides, the bond angles between nucleotide triplets, and the torsion angles, as shown in Figure S3. It can be observed that the structures of the RNA we generated are similar in distribution to those of real RNA, capable of reproducing the characteristics of these local structures.
B.3 Experimental Results on Few-shot Set
@@
Following the experimental setup of Section 5.2, we conduct similar experiments on the few-shot set, with the experimental results as shown in Table S2. It can be observed that despite the model being exposed to only a small amount of sample information during training, it still demonstrated excellent performance. In addition, we have also presented the detailed results of vina and pose scores for these fifteen RNA-ligand pairs in the Figure S8, fully demonstrating the experimental performance of different model variants.
B.4 Using RiboFlow for De Novo RNA Design
@@
Pipeline. For ligand A2F. We employ RiboFlow to generate RNA sequences of length 67, which are then docked against A2F ligand-bound conformers using GNINA. A specific RNA candidate is selected based on a composite score integrating Vina and ligand pose scores. The corresponding ground-truth RNA sequence and the designed RNA sequence are presented in the results.
For ligand GNG. We first leverage RiboFlow to generate RNAs ranging from 50 to 150 nucleotides in length. These RNA sequences are then docked against GNG conformers, generated by RDKit, using GNINA. A 90-nucleotide RNA candidate is selected based on a composite score that integrates vina and ligand pose scores. To validate this candidate, gRNAde is applied to determines its corresponding sequence, and AlphaFold3 predicts the structure of the resulting RNA-GNG complex.
@@
For ligand GNG. The AlphaFold3-predicted complex structure exhibits a TM-score of 0.41 and an RMSD of 3.82 Å compared to the RiboFlow-designed RNA backbone, demonstrating the validity of the designed RNA. Furthermore, as illustrated in Figure S4, a high consistency is observed between the GNINA-predicted binding pocket and the AlphaFold3-predicted binding region. These findings provide strong support for RiboFlow’s potential to facilitate the design of RNAs with high binding affinity.
Figure S4: De novo design of a 90-nucleotide RNA corresponding to the ligand GNG with a molecular weight of 267.2 Da. The RNA-ligand complex structure is predicted by AlphaFold3, whereas another is generated by RiboFlow and docking using GNINA. The AlphaFold3-predicted complex structure exhibits a TM-score of 0.41 and an RMSD of 3.82 Å compared to the RiboFlow-designed RNA backbone, demonstrating the validity of the designed RNA.
Table S2: Comparison of model performance on the few-shot evaluation set. The best results are highlighted in bold, while the second-best results are underlined. The table presents two experimental results: one based on a fixed RNA length and ligand-bound structure, and another based on ligand-bound structures with RNA length sampling. Colored boxes denote different experimental setups.
Vina. ()($\downarrow$)( ↓ ) Affinity ()($\uparrow$)( ↑ ) %Pose. ()($\uparrow$)( ↑ ) AF. ()($\uparrow$)( ↑ ) %Val.()($\uparrow$)( ↑ ) Div.()($\uparrow$)( ↑ ) Nov.()($\downarrow$)( ↓ )
Top10 Median Top10 Median
Generation with Given RNA Length and Ligand Conformation
Random -2.16 -2.03 2.89 2.72 5.10 7.11 - - -
RNA-FrameFlow -3.95 -3.88 3.28 3.11 52.2 16.4 20.3 0.279 0.570
RiboFlow -4.63 -4.49 3.68 3.55 67.2 22.8 9.68 0.216 0.645
+ Pre -4.74 -4.63 4.06 3.82 76.0 47.1 32.4 0.310 0.534
+ Crop -4.56 -4.51 4.08 3.81 73.3 25.6 18.1 0.257 0.623
+ Pre + Crop -4.47 -4.41 4.05 3.79 78.7 49.9 25.9 0.331 0.527
Generation with Given RNA Length and Ligand Conformation
Random -2.23 -2.15 3.09 2.93 7.00 8.84 - - -
RNA-FrameFlow -4.35 -4.01 3.31 3.16 55.4 19.1 22.0 0.521 0.580
RiboFlow -4.87 -4.66 3.90 3.78 70.4 24.1 10.2 0.529 0.664
+ Pre -5.04 -4.89 4.26 4.13 73.8 48.6 35.7 0.572 0.520
+ Crop -4.91 -4.75 4.28 4.06 72.0 32.4 17.4 0.535 0.622
+ Pre + Crop -4.96 -4.82 4.21 4.11 76.2 51.3 26.2 0.577 0.539
Appendix C Flow Matching
Flow Matching (FM) (Lipman et al., 2023) is a method designed to reduce the simulation required for learning Continuous Normalizing Flows (CNFs), which is a class of deep generative models that generate data by integrating an ordinary differential equation (ODE) over a learned vector field. In this section, we will provide a concise overview of the flow matching approach.
$\frac{d}{dt}\phi_{t}(x)=v_{t}(\phi_{t}(x)),\quad\phi_{0}(x)=x.$ (17)
Once trained, samples can be generated by simulating Equation 17 using the learned vector field $v_{t}$.
Appendix D Implementation Details
RiboFlow utilizes several hyperparameters in the experiments, which are crucial for the model’s training and sampling processes. Therefore, we provide some key hyperparameters to facilitate the reproduction of our experiments. The optimal hyperparameters are indicated in bold.
Table S3: Hyperparameters for the RiboFlow.

### v3

@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This study has been supported by the National Natural Science Foundation of China [62041209], Natural Science Foundation of Shanghai [24ZR1440600], the Science and Technology Commission of Shanghai Municipality [24510714300], the Lingang Lab Fund [LGL-8888].
APPENDIX
7 Dataset
7.1 Dataset Preparation
@@
(i) Data collection. We obtain raw structural data in mmCIF format from the RCSB PDB database using the advanced search query, which supports custom filtering based on specific structural attributes. The search is restricted to structures containing at least one RNA polymer chain and one ligand instance. These criteria are defined using the following query settings: Structure Attributes -> Number of Distinct RNA Entities >= 1 and Structure Attributes -> Total Number of Non-polymer Instances >= 1. This step serves as the initial data collection phase.
(ii) Ligand filtering. We adopt a portion of the ligand selection criteria from the HairBoss database [33], while relaxing certain rigid druggability restrictions, aiming to balance ligand diversity and effective identification. Specifically, the following rules are applied during our workflow: (1) the molecular weight of the ligand must fall within the range of 100-1000 Daltons, and (2) the ligand must not be a solvent molecule or a metal ion. Consequently, we expand the original ligand library beyond the 311 ligand classes used in HairBoss by incorporating an additional 237 ligand classes, resulting in a total of 548 unique ligand classes.
@@
(iv) Redundancy reduction. To reduce redundancy and improve dataset quality, we employ MMseqs2 [36] to cluster RNA sequences at a 90% sequence identity threshold. Notably, RNA sequences within the same cluster but bound to different ligands are still treated as distinct RNA-ligand pairs.
Figure 4: The collection pipeline of RiboBind dataset.
Figure 5: An example of RNA dynamic cropping. Based on a selected nucleotide, we randomly select a cropping length within the range of 30 to 200 to obtain the trimmed RNA-ligand pair for amplifying the existing data.
Through this comprehensive pipeline, we construct the RiboBind dataset, offering a substantially larger and more diverse collection of RNA-small molecule interactions compared to existing datasets. The final RiboBind dataset includes 1,591 RNA-ligand complexes and 3,012 RNA-ligand pairs. In comparison, as of December 2024, the publicly accessible HairBoss dataset comprises only 862 RNA-ligand complexes and 1,471 RNA-ligand pairs. This notable increase in both scale and diversity highlights the value of RiboBind as a robust resource for RNA-small molecule interaction studies. For more detailed statistical information about RiboBind, please refer to Figure 6.
7.2 Dataset Comparison
Currently, there are also some efforts to collect and curate large-scale RNA-ligand datasets. In addition to the HariBoss dataset referred to in this article, some well-known datasets include PDBBind [30] and RNAmigos2 [8]. PDBBind is a comprehensive database primarily focused on experimentally measured binding affinity data for protein–ligand complexes, though it also includes a small subset of RNA–ligand complexes. In its latest version (v.2024), it contains 234 RNA–ligand complexes. According to our inspection, all of these entries are already included in the RiboBind dataset, making PDBBind essentially a high-quality subset of RiboBind in the context of RNA–ligand interactions.
The RNAmigos2 dataset, on the other hand, was introduced alongside the recently proposed RNA structure-based virtual screening tool RNAmigos2. This dataset includes 1,740 experimentally validated RNA–ligand binding site structures. After deduplication, we verified that all of these experimental structures are also present in the RiboBind dataset. In addition to the experimental data, the RNAmigos2 dataset also contains 1.3 million synthetic affinity data points, generated via in silico molecular docking. These synthetic interactions were computed by docking approximately 800 ligands (sourced from the ChEMBL database) against the 1,740 experimentally solved RNA structures in all pairwise combinations. Since RiboBind currently only includes experimentally determined structural data, it remains fully consistent with the experimental portion of the RNAmigos2 dataset. However, we plan to incorporate the synthetic data in future versions of RiboBind to increase the chemical diversity and overall scale of the dataset, thereby enhancing the generative performance of our models.
7.3 Data Augmentation
@@
In order to ensure that the cropped RNA sequences retain interactions with small molecules, the cropping strategy is informed by interaction data from existing datasets. Specifically, we select RNA bases involved in interactions with the small molecules as candidate bases. By calculating the total interaction counts between each base and every atom in the small molecules, we can identify the top three bases with the highest interaction frequencies as cropping centers. The cropping length is randomly sampled within the range of 30–200 nucleotides. Using the selected base as the center, the RNA sequence is cropped symmetrically, with half the total length taken from both upstream and downstream of the base, forming the final RNA-ligand pair. For more detailed statistical information about the amplified RiboBind dataset, please refer to Figure 7. This approach ensures that the cropped RNA sequences retain critical interaction information for downstream tasks.
@@
(a) Distribution of RNA Length
(b) Distribution of Ligand Type
Figure 6: Statistics of the originally collected RiboBind dataset. (a) The number of RNAs within different length ranges, and the vast majority of RNAs exceed the 200-nucleotide limit; (b) The proportion of the top ten categories of ligands.
(a) Distribution of RNA Length
(b) Distribution of Ligand Type
Figure 7: Statistical analysis of the RiboBind dataset, amplified using a dynamic cropping strategy after removing sequences shorter than 30 nucleotides, reveals the following: (a) The amplified dataset exhibits a more balanced distribution of RNA lengths compared to the original RiboBind dataset, making it better suited for model training. (b) The top ten ligand categories in the amplified dataset are still predominantly composed of the seven major ligands (FME, GTP, SAM, PHE, SPM, MPD, and SPD) found in the original dataset. This suggests that the ligand bias introduced by the amplification process remains within an acceptable range.
8 More Results and Analysis
8.1 The Impact of Pre-training Steps on RNA Generation
To fully demonstrate the impact of different pre-training steps on RNA generation, we comprehensively show the changes in model performance under various pre-training steps. We additionally introduced the scRMSD metric to evaluate the performance changes. scRMSD refers to self-consistency Root Mean Squared Deviation between the generated and the predicted backbone atoms to reflect structural validity. During the pre-training stage, a batch size of 32 is selected, utilizing 4 A100 80GB accelerator cards, completing 200K pre-training steps in nearly 20 hours. The experimental results are shown in Table 4. 50K_step_50 means that the model samples with a sampling step of 50 after training for 50K steps.
Table 4: Detailed results of unconditional generation under different pre-training steps.
@@ 8.1 The Impact of Pre-training Steps on RNA Generation
Figure 8: Comparison of probability distribution histograms for nucleotide bond distances, bond angles, and torsion angles between generated and real RNA structures. Top: RNA-FrameFlow-R; Bottom: RiboFlow. RiboFlow shows a closer match to the real data distribution, particularly in bond angles, indicating improved geometric realism over RNA-FrameFlow-R.
8.2 The Local Structure of RNA Analysis
@@
To demonstrate the validity of the generated RNA structures, we conduct a detailed comparison between the RNA structures generated by RNAFrame-Flow-R and RiboFlow within the [50,150] range and the actual RNA structure distributions. We primarily illustrate histograms of the probability distributions for the bond distances between nucleotides, the bond angles between nucleotide triplets, and the torsion angles, as shown in Figure 8. It can be observed that the structures of the RNA we generated are similar in distribution to those of real RNA, capable of reproducing the characteristics of these local structures.
8.3 The Impact of Sampling Lengths on Model Generation Quality
Due to the pre-training data distribution limitations in RNASolo (which is primarily concentrated in the range of 30–200), we set the RNA sampling range to [50, 150]. However, we are willing to explore RiboFlow’s performance in designing longer RNA sequences. To this end, we conducted additional experiments, performing unconditional RNA backbone design for sequences of lengths [200, 250, 300, 350, 400, 450]. Notably, attempting longer sequences on an RTX 4090 GPU (24GB memory) triggers an OOM error due to the memory requirements of RhoFold. The scTM experimental results for each length are as shown in Figure 9.
Figure 9: The impact of sampling lengths on model generation quality.
Our findings indicate that for longer sequences, RNA validity declines significantly compared to the performance in the [50, 150] range reported in the paper. However, it is important to note that RhoFold explicitly states that its training data only includes nucleotide sequences of lengths 16–256, and its accuracy beyond this range has not been fully validated. Therefore, this experiment provides only a rough assessment of the model’s generalization ability. We look forward to advancements in RNA engineering, where more powerful structural prediction models will enable more complex RNA designs.
8.4 Using Different Backbone Atoms for Model Evaluation
In the paper, we used the $C4^{\prime}$ atom to evaluate model performance in experimental results. To assess the validity of generated RNA structures, we employed scTM and scRMSD metrics. The TM-score was calculated using US-align [49], which by default selects a single backbone atom for structural evaluation.
Recognizing that a multi-atom approach could provide a more comprehensive assessment, we expand the evaluation to include the backbone atoms available in US-align ($C3^{\prime}$, $C4^{\prime}$, $C5^{\prime}$, see https://zhanggroup.org/US-align/ for detailed information). We computed scTM and scRMSD for each individual atom and their mean values to enhance evaluation robustness. Note that scRMSD refers to the RMSD between the structure predicted by RhoFold and the structure generated by the model. We refer to the settings in RNA-FrameFlow [3] and define samples with scRMSD below 4.3 Å as valid structures. We report the performance of the model using different backbone atoms under unconditional generation tasks.
The results can be found in Table 5. Please note that each column in the figure represents the validity of the structure generated by using different metrics for calculation. For example, %scRMSD(Avg.) represents the proportion of valid structures generated by calculating scRMSD using the average metric of these three atoms $C3^{\prime}$, $C4^{\prime}$, and $C5^{\prime}$ simultaneously. It can be indicate that evaluations based on multiple atoms yield consistent performance trends, supporting a more accurate reflection of the model’s structural generation quality.
Table 5: Structure generation validity of the model under unconditional generation, evaluated using different backbone atoms ($C3$’, $C4$’, $C5$’) and their average.
Method %scRMSD(Avg.) %scRMSD($C3^{\prime}$) %scRMSD($C4^{\prime}$) %scRMSD($C5^{\prime}$) %scTM(Avg.) %scTM($C3^{\prime}$) %scTM($C4^{\prime}$) %scTM($C5^{\prime}$)
RNA-FrameFlow
Sample_Step_50 27.0 27.7 27.0 27.0 26.3 27.0 26.7 26.7
Sample_Step_100 29.3 29.7 30.3 29.7 31.6 32.0 31.3 30.6
RNA-FrameFlow-R
Sample_Step_50 25.7 25.7 25.0 25.7 26.7 26.0 26.0 25.7
Sample_Step_100 26.3 26.3 26.0 26.7 30.0 31.3 30.0 31.6
RiboFlow-Codesign
Sample_Step_50 32.7 32.3 32.7 32.3 36.7 38.3 37.0 36.3
Sample_Step_100 31.3 31.0 31.3 31.3 33.7 35.7 33.3 31.3
Table 6: Comparison of RNA generation by structure-based evaluation. The best results are highlighted in bold, while the second-best results are underlined. The table presents two experimental results: one based on a fixed RNA length and ligand-bound structure, and another based on ligand-bound structures with RNA length sampling. Colored boxes denote different experimental setups.
Vina. $(\downarrow)$ GerNA. $(\uparrow)$ %AF. $(\uparrow)$
Top10 Median
Generation Given Ligand Structure and True Length
Random -2.98 -2.85 0.121 5.32
RNA-FrameFlow -4.33 -4.25 0.225 23.9
RiboFlow -5.07 -5.01 0.231 25.2
+ Pre -5.32 -5.26 0.452 52.3
+ Crop -5.17 -5.09 0.389 40.2
+ Pre + Crop -5.43 -5.39 0.439 49.7
Generation with Given RNA Length and Ligand Conformation
Random -2.32 -2.15 0.110 6.47
RNA-FrameFlow -4.35 -4.24 0.217 17.4
RiboFlow -5.36 -5.21 0.245 23.1
+ Pre -5.64 -5.49 0.437 45.3
+ Crop -5.41 -5.30 0.402 33.5
+ Pre + Crop -5.56 -5.38 0.393 41.3
8.5 Experimental Results on Structure-based Evaluation
Following the experimental setup of Section 5.2, we conduct similar experiments on the structure-based evaluation set, with the experimental results as shown in Table 6. Our findings show that RiboFlow and its variants still exhibit strong performance under structural-based division.
8.6 Experimental Results on Few-shot Evaluation
@@
Following the experimental setup of Section 5.2, we conduct similar experiments on the few-shot set, with the experimental results as shown in Table 7. It can be observed that despite the model being exposed to only a small amount of sample information during training, it still demonstrated excellent performance. In addition, we have also presented the detailed results of vina scores for these fifteen RNA-ligand pairs in the Figure 13, fully demonstrating the experimental performance of different model variants.
Table 7: Comparison of model performance on the few-shot evaluation set. The best results are highlighted in bold, while the second-best results are underlined. The table presents two experimental results: one based on a fixed RNA length and ligand-bound structure, and another based on ligand-bound structures with RNA length sampling. Colored boxes denote different experimental setups.
Vina. $(\downarrow)$ GerNA. $(\uparrow)$ AF. $(\uparrow)$
Top10 Median
Generation with Given RNA Length and Ligand Conformation
Random -2.16 -2.03 0.125 7.11
RNA-FrameFlow -3.95 -3.88 0.206 16.4
RiboFlow -4.63 -4.49 0.276 22.8
+ Pre -4.74 -4.63 0.415 47.1
+ Crop -4.56 -4.51 0.401 25.6
+ Pre + Crop -4.47 -4.41 0.376 49.9
Generation with Given RNA Length and Ligand Conformation
Random -2.23 -2.15 0.110 8.84
RNA-FrameFlow -4.35 -4.01 0.207 19.1
RiboFlow -4.87 -4.66 0.287 24.1
+ Pre -5.04 -4.89 0.433 48.6
+ Crop -4.91 -4.75 0.412 32.4
+ Pre + Crop -4.96 -4.82 0.403 51.3
8.7 Sampling Time Efficiency Analysis
Figure 10: Sampling time comparison between RiboFlow and RNA-FrameFlow. Time in seconds is plotted against RNA sequence length (nt) for both 50 and 100 sampling steps.
In this section, we evaluate the sampling efficiency of RiboFlow in comparison to RNA-FrameFlow, a critical factor for the practical utility of generative models. This analysis measures the time required to generate RNA structures of varying lengths (approximately 50 to 300 nucleotides) using both 50 and 100 sampling steps. All experiments were conducted on an identical hardware platform to ensure a fair and consistent comparison of computational overhead.
Figure 10 illustrates the sampling time costs. At 50 sampling steps, RiboFlow exhibits a sampling time that is largely comparable to RNA-FrameFlow across the tested sequence lengths, with RiboFlow’s time being slightly higher in most instances. When the number of sampling steps is increased to 100, RiboFlow incurs a more noticeable increase in computational time relative to RNA-FrameFlow. This difference in sampling time between the two models becomes more pronounced with increasing sequence length at 100 steps, indicating that the processing of additional conditioning information in RiboFlow contributes to a steeper scaling of its sampling time, particularly for longer sequences. Nevertheless, the sampling times for RiboFlow remain within a practical and acceptable range for typical applications.
8.8 Using RiboFlow for De Novo RNA Design
@@
Pipeline. For ligand A2F. We employ RiboFlow to generate RNA sequences of length 67, which are then docked against A2F ligand-bound conformers using AutoDock Vina. A specific RNA candidate is selected based on a composite score integrating Vina scores and ligand pose validity. The corresponding ground-truth RNA sequence and the designed RNA sequence are presented in the results.
For ligand GNG. We first leverage RiboFlow to generate RNAs ranging from 50 to 150 nucleotides in length. These RNA sequences are then docked against GNG conformers, generated by RDKit, using AutoDock Vina. A 90-nucleotide RNA candidate is selected based on a composite score that integrates vina and ligand pose validity. To validate this candidate, gRNAde is applied to determines its corresponding sequence, and AlphaFold3 predicts the structure of the resulting RNA-GNG complex.
@@
For ligand GNG. The AlphaFold3-predicted complex structure exhibits a TM-score of 0.41 compared to the RiboFlow-designed RNA backbone, demonstrating the validity of the designed RNA. Furthermore, as illustrated in Figure 11, a high consistency is observed between the GNINA-predicted binding pocket and the AlphaFold3-predicted binding region. These findings provide strong support for RiboFlow’s potential to facilitate the design of RNAs with high binding affinity.
Figure 11: De novo design of a 90-nucleotide RNA corresponding to the ligand GNG with a molecular weight of 267.2 Da. The RNA-ligand complex structure is predicted by AlphaFold3, whereas another is generated by RiboFlow and docking using GNINA. The AlphaFold3-predicted complex structure exhibits a TM-score of 0.41 compared to the RiboFlow-designed RNA backbone, demonstrating the validity of the designed RNA.
9 More Details of RiboFlow Training and Generation
9.1 Model Training Details
RiboFlow utilizes several hyperparameters in the experiments, which are crucial for the model’s training and sampling processes. Therefore, we provide some key hyperparameters to facilitate the reproduction of our experiments in Table 8. The optimal hyperparameters are indicated in bold.
Table 8: Hyperparameters for the RiboFlow.